OpenStudy (anonymous):

Use the functions to answer the question. f(x) = 4x – 1; g(x) = 5x^2 Which is equal to (f * g) (x)? 80x^2 – 40x + 5 20x^3 – 5x^2 5x^2 + 4x – 1 20x^2 – 1

OpenStudy (anonymous):

This is obviously composite functions, so (f º g)(x) = f(g(x)) = 4x - 1 = 4(5x^2) - 1 = 20x^2 - 1. So, the answer is D, the last one.

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (anonymous):

You welcome, a medal would be nice if you are satisfied :).

OpenStudy (anonymous):

I already put your answer as the best response.

OpenStudy (anonymous):

Okay, thanks, :). Good luck with your homework/test/etc.

OpenStudy (anonymous):

Thank you!